# Developer Perspectives on Licensing and Copyright Issues Arising from Generative AI for Software Development

### v3

@@ Abstract.
Despite the utility that Generative AI (GenAI) tools provide for tasks such as writing code, the use of these tools raises important legal questions and potential risks, particularly those associated with copyright law. As lawmakers and regulators engage with those questions, the views of users can provide relevant perspectives. In this paper, we provide: (1) a survey of 574 developers on the licensing and copyright aspects of GenAI for coding, as well as follow-up interviews; (2) a snapshot of developers’ views at a time when GenAI and perceptions of it are rapidly evolving; and (3) an analysis of developers’ views, yielding insights and recommendations that can inform future regulatory decisions in this evolving field. Our results show the benefits developers derive from GenAI, how they view the use of AI-generated code as similar to using other existing code, the varied opinions they have on who should own or be compensated for such code, that they are concerned about data leakage via GenAI, and much more, providing organizations and policymakers with valuable insights into how the technology is being used and what concerns stakeholders would like to see addressed.
@@ 1. Introduction
Generative AI (GenAI) tools have become widely adopted across many different domains, including software engineering (SE)  (R. Li et al., 2023; Ebert and Louridas, 2023; Rajbhoj et al., 2024; Krill, [n. d.]). Several GenAI coding assistants, including GitHub’s Copilot (GitHub, [n. d.]), Tabnine (Tabnine, [n. d.]), Codeium (codeium, [n. d.]), and Cody (cody, [n. d.]), as well as general purpose tools such as ChatGPT (OpenAI, [n. d.]), Claude (Anthropic Models, [n. d.]), and Gemini (Gemini Team et al., 2023), have become readily accessible, either as IDE extensions or standalone applications, enabling developers to perform many coding tasks with little effort, including automated code completion, summarization, and debugging. A 2024 Stack Overflow survey reported that 76 percent of 65,000 participating developers were using or planning on using AI coding tools (May, [n. d.]); more than one million developers used Copilot (GitHub, [n. d.]) in its first year (Yao, 2023). The popularity of these tools stems from the fact that they are easy to access and can generate diverse content with relatively little effort, thus helping users perform many tasks more efficiently (Liang et al., 2024; Barke et al., 2023). Specifically, these emerging GenAI tools can help developers with various software engineering tasks, including code autocompletion, summarization, and debugging.
@@
However, in addition to these benefits, the use of GenAI tools can introduce various issues, including bias or discrimination (Ferrara, 2023), security threats (Majdinasab et al., 2024a), and compromise of private information (Gupta et al., 2023). GenAI can also pose legal risks related to intellectual property concerns, including copyright infringement (Generative Artificial Intelligence and Copyright Law, [n. d.]). Many legal questions regarding the use of GenAI tools remain unanswered, and any answers will likely vary among jurisdictions and across cases.
@@
As of this writing, there are several pending lawsuits against high-profile providers of GenAI tools including OpenAI (Authors Guild v. OpenAI Inc., 2023), StabilityAI (get, 2023), and Midjourney (Andersen v. Stability AI Ltd., [n. d.]), as well as active consideration by the U.S. Copyright Office and other governmental entities around the world of the copyright issues raised by the use of GenAI (Office, 2024a, 2023; Council of the European Union, 2024b; Office, 2025). These matters implicate several considerations, including the extent to which works associated with the use of GenAI, including prompts and models, are protected by copyright and, if so, who owns the rights to such works; whether the use of open source software and other protected works as training data results in license violations or otherwise constitutes copyright infringement; and whether output resulting from the use of GenAI that is similar to preexisting work constitutes copyright infringement.
@@
Because any regulation of this space– whether by governmental entities or corporate entities– should take account of current and future practices, we aim in this paper to identify developers’ views on (1) using GenAI tools for software development tasks and (2) associated legal concerns, focusing specifically on copyright law. This is motivated by recent solicitations from the United States Copyright Office requesting the opinions of experts, organizations, and individuals (e.g., developers) for active consideration in the development of future regulation (Office, 23).
@@
This paper presents a study conducted by a joint team of SE and legal researchers that surveyed 574 software developers worldwide who use GenAI tools for software development tasks (particularly code generation). Through an online survey and follow-up interviews, we investigated developers’ opinions on potential emerging legal issues, the perception of what is copyrightable, ownership of generated code, and related considerations. We also sought to understand potential developer misconceptions, assess the impact of GenAI on their work, and assess their awareness of licensing/copyright risks.
@@
Using qualitative and quantitative methods to analyze survey responses, we found that developer opinions on copyright issues vary broadly, particularly on the topic of model output ownership, and that many developers are aware of the nuances and complications associated with answering these complex legal questions. We discuss the results of our study, interpret the findings under the background of U.S. law, and offer insights for future work. We focus on developers rather than other stakeholders, such as compliance officers and policymakers, because developers are often the ones who directly interact with GenAI during software development. We survey an international population of developers because software development is a global activity, particularly for open-source projects. Additionally, large technology companies regularly employ individuals from various countries (and jurisdictions) around the world. Given this, it is appropriate to consider the thoughts, opinions, and perspectives of developers regardless of their geographic location.
@@
We make available our survey, results, and other artifacts for transparency and validation in an online replication package (ano, 2024).
@@ 2. Background and Related Work
This intersectional field of copyright and generative AI relies on concepts drawn both from the legal domain and from software engineering. We discuss these concepts in this section as well as prior work related to our study.
@@ 2.1. Legal background
Unless it is in the public domain, computer code is generally protected by copyright in the U.S. so long as it constitutes an original work of authorship and is fixed in a tangible medium of expression, such as on a server (17 U.S.C. § 102, a). Similar provisions exist in the intellectual property legislation of other countries, including those in the European Union (Commission, 2024). The owner of a copyright—which may be one or more individuals or corporate entities—has several exclusive rights under U.S. copyright law, including the rights to reproduce the work in copies, to create derivative works (works based on the copyrighted work), and to distribute the work, which also includes the right to authorize others to engage in such activities (17 U.S.C. § 106, [n. d.]). In addition, U.S. copyright law prohibits the “removal or alteration of copyright management information,” defined as including “the name of, and other identifying information about, the author of a work,” although the limits of this provision and its application to GenAI are as yet uncertain (17 U.S.C. § 1202, b).
The use of GenAI implicates these rights in several ways that are currently under consideration by the U.S. Copyright Office and the courts. (Office, 2023; The New York Times Company v. Microsoft Corporation, [n. d.]; Brittain, 2023; cop, 2023). For example, the plaintiffs in Doe v. GitHub (DOE 1 et al v. GitHub, [n. d.]), filed in November 2022, allege that Copilot removes copyright management information and violates the copyright licenses of the plaintiffs’ open source software by using that code as training data and generating code that is a near-identical reproduction of the plaintiffs’ code but without adhering to the terms of the associated licenses.
@@
Legislation both within and outside the U.S. also has implications for GenAI and copyright. Various U.S. states have attempted to enact legislation to govern GenAI (colorado-bill, 2024; california-regulation, 2024). In March 2024, the European Parliament approved the EU AI Act (Council of the European Union, 2024b), which, among other things, requires the providers of general-purpose AI models to “put in place a policy to comply with Union law on copyright and related rights, and in particular to identify and comply with, including through state-of-the-art technologies, a reservation of rights expressed pursuant to Article 4(3) of Directive (EU) 2019/790” (Council of the European Union, 2024a). This requires that creators of general-purpose AI models ensure upstream copyright compliance—apply due diligence in using copyrighted materials as part of their training data, obtain consent from copyright owners, and allow copyright holders to opt out—and ensure downstream copyright compliance, mitigating the risk—often due to overfitting—of generating verbatim content that is too similar to the (copyrighted) original. In 2023, an early announcement of the EU Global Principles for Artificial Intelligence, which led to the EU AI Act, encouraged some organizations, including French media organizations France Médias Monde and TF1, to forbid companies such as OpenAI from mining their data (Karathanasis, 2023); that year also saw a temporary ban on ChatGPT in Italy (McCallum, 2023). Even countries like Japan, which initially affirmed that training AI models on protected content was not a violation of copyright law (Lanz, [n. d.]), have started to consider additional restrictions designed to protect copyright holders (Shinbun, 2023). Gervais et al. (Gervais et al., 2024) provide an in-depth overview of the then-current copyright landscape around the globe.
@@ 2.2. Technical background
Additional work relevant to GenAI and copyright issues focuses on research that may help answer some of the technological questions raised above. For example, techniques exist to potentially reveal whether a machine learning model was trained on a certain data record, such as membership inference attacks (Duan et al., 2024; Hu et al., 2022; Majdinasab et al., 2024b; Zhang and Li, 2023; Z. Yang et al., 2023). Such attacks can extract training data from the target model’s output, including private and personal information (Carlini et al., 2021). Memorization of training data is a known issue, including in models trained on code (Yang et al., 2024), with research suggesting that typical measures aimed at preventing verbatim extraction of training data are not sufficient to prevent the data from being obtained with other methods, such as by utilizing style-transfer prompts (Ippolito et al., 2023). Tools like CopyrightCatcher (CopyrightCatcher, [n. d.]) have been developed to detect the presence of copyrighted material in model output, evidencing that this is an issue of concern to developers.
@@ 2.3. Related work
Prior research has explored several questions and concerns regarding copyright and GenAI (Lucchi, 2023; Guadamuz, 2024). For example, Lee et al. (Lee et al., 2023a) outline the intricate GenAI supply chain and its intersection at various points with aspects of U.S. copyright law, revealing that answers to copyright questions may depend on the specific details of how an individual model was trained. Craig (Craig, 2024) has cautioned against applying or expanding copyright principles to address issues raised by GenAI without first carefully considering the unintended consequences of such an approach. Kugler (Kugler, 2024) provides a brief overview of the copyright challenges posed by GenAI and how various governments around the globe have begun addressing them.
Other research has considered the copyright implications of the emerging base model and subsequent fine-tuning paradigm (Yew, 2024) and has studied the applicability of fair use for foundation models (Henderson et al., 2023). Additional research (Neel and Chang, 2023) presents a review of the discourse surrounding copyright and privacy issues related to LLMs. The research most closely related to our study is by Liang et al. (Liang et al., 2024), who investigated why developers choose to use AI programming assistants and the challenges they encountered.
Our work builds on this related work by contributing an in-depth exploration of developers’ perceptions of the copyright issues surrounding the use of AI-based code-generation tools, both in the U.S. and elsewhere around the world. Understanding these perspectives allows future regulation in this area to be informed by beliefs on the ground rather than in a vacuum. The U.S. Copyright Office, to “inform the Office’s study and help assess whether legislative or regulatory steps in this area are warranted,” has actively sought public comment on copyright issues pertaining to artificial intelligence (Office, 23). Even after receiving over 10,000 comments, the Copyright Office continues to study this developing area (Office, 2024b, 2025). In the same spirit, our work aims to capture the practices and views of developers to provide a resource for policymakers.
@@ 3. Study Methodology
The goal of this study is to investigate developers’ perceptions about licensing and copyright issues related to the use of GenAI technology for software development. (Although we use the term “software development” throughout the paper, our study primarily focuses on the use of GenAI for code generation as the main task that can lead to copyright and licensing issues.) The primary quality focus is to provide lawmakers, regulators, and other interested parties with a summary and analysis of relevant developer perspectives pertaining to the copyright and licensing questions that arise when developing software with the assistance of GenAI. The context consists of 574 developers of open-source projects hosted on GitHub, whom we surveyed via an online survey and follow-up interviews.
@@
    What other legal concerns do developers anticipate as the use of GenAI increases? There are a plethora of other potential legal issues to consider as GenAI becomes more widely adopted, including data privacy, generation of malicious content, and tort liability. In this RQ, we report developers’ perspectives on these topics which spread beyond copyright law.
@@ 3.1. Survey design
To design our survey, we followed general guidelines (Groves et al., 2009) as well as SE-specific best practices (Pfleeger and Kitchenham, 2001; Kitchenham and Pfleeger, 2002a, b, c, d, 2003). The final questionnaire went through multiple iterations of review and improvements from six SE researchers and one law researcher, all authors of this paper. All researchers contributed to the design of the study, but in particular, the SE researchers contributed the domain experience to ask questions about procedures and practices relevant to developers, and the law researcher facilitated the identification of key areas that could be of concern related to copyright, licensing, etc., allowing the survey to ask the most pertinent questions possible. We carefully formulated the questions of the survey and ensured they were written clearly and concisely to avoid biasing respondents. We also conducted a pilot study with graduate students from our research lab who actively use GenAI tools. Based on their feedback, we further refined our questionnaire, improving questions that were ambiguous or hard to understand.
@@
  • The section on “Current use of GenAI for coding” (12 questions) asked about respondents’ experience using GenAI tools for code generation (C1, C2, C8), including the tools they used (C3); the development tasks for which they used the tools (C4); the benefits and challenges of using such tools (C5, C6); whether one examines the Terms of Service for tools (C7); organizational policies of using such tools (C9, C12); and procedures to document tool usage (C10, C11).
  • The section on “Understanding and Perception of Copyright Issues” (14 questions) asked about developers’ perceptions of copyright issues surrounding GenAI, including their views on if and how copyright, litigation, and legal risks factor into the use of GenAI (U1, U2, U3); the effects of these concerns on developers’ workflows (U4); whether the use of source code in training models should require attribution and/or monetary compensation (U6, U5); the use of their code in model training without permission (U7); having models produce code similar to theirs (U8); reusing without permission a prompt that they created (U9); ownership and copyrightability of generated code (U11, U12, U13); and other, broader concerns about copyright/legal issues stemming from development tasks with AI-based tools (U14).
  • The “Demographics” section (8 questions) asked about respondents’ age (D1); experience with SE and AI (D2, D3); their primary programming languages (D4); whether they develop open source or proprietary software (D5); their location (D6); and if they have received training on copyright (D6, D7).
The survey was designed to be completed in about 15 to 20 minutes and consisted mostly of multiple-choice questions (26), complemented by open-ended questions (7) that allowed participants to elaborate on their responses. Because we included logic in our survey to ask only those questions relevant to a participant’s indicated experience, not every participant saw every question. We included this logic to avoid asking questions about topics the participants indicated they did not have experience with. For example, if a participant indicated that they were not familiar with the practice of generating source code using GenAI (C1), we did not ask them questions about generating source code (C2-C12). The complete survey text and a brief description are found in our replication package (ano, 2024).

### v5

@@ Abstract.
Despite the utility that Generative AI (GenAI) tools provide for tasks such as writing code, the use of these tools raises important legal questions and potential risks, particularly those associated with copyright law. As lawmakers and regulators respond to these questions, the views of users can offer relevant perspectives. In this paper, we provide: (1) a survey of 574 developers on the licensing and copyright aspects of GenAI for coding, as well as follow-up interviews; (2) a snapshot of developers’ views at a time when GenAI and perceptions of it were rapidly evolving; and (3) an analysis of developers’ perspectives, yielding insights and recommendations that can inform future regulatory decisions in this evolving field. Our results show the benefits developers derive from GenAI, how they view the use of AI-generated code as similar to using other existing code, the varied opinions they have on who should own or be compensated for such code, that they are concerned about data leakage via GenAI, and other findings, providing organizations and policymakers with valuable insights into how the technology is being used and the concerns that stakeholders believe warrant attention.
open-source software, generative ai, machine learning, large language models, qualitative research
copyright: ccjournal: TOSEMjournalyear: 2025journalvolume: 1journalnumber: 1article: 1publicationmonth: 1doi: 10.1145/3743133ccs: Social and professional topics Copyrightsccs: Social and professional topics Licensingccs: Computing methodologies Artificial intelligence
The first two authors contributed equally to this work.
@@ 1. Introduction
Generative AI (GenAI) tools have been widely adopted in many different domains, including software engineering (SE)  (R. Li et al., 2023; Ebert and Louridas, 2023; Rajbhoj et al., 2024; Krill, [n. d.]). Several GenAI coding assistants, including GitHub Copilot (GitHub, [n. d.]), Tabnine (Tabnine, [n. d.]), Codeium (codeium, [n. d.]), and Cody (cody, [n. d.]), as well as general-purpose tools such as ChatGPT (OpenAI, [n. d.]), Claude (Anthropic Models, [n. d.]), and Gemini (Gemini Team et al., 2023), have become readily accessible, either as IDE extensions or stand-alone applications, enabling developers to perform many coding tasks, including automated code completion, summarization, and debugging. A 2024 Stack Overflow survey reported that 76 percent of 65,000 participating developers were using or planning to use AI coding tools (May, [n. d.]); more than one million developers used Copilot (GitHub, [n. d.]) in its first year (Yao, 2023). The popularity of these tools comes from the fact that they are easy to access and can generate various types of content with relatively little effort, thus helping users perform many tasks more efficiently (Liang et al., 2024; Barke et al., 2023).
@@
In addition to these benefits, however, the use of GenAI tools can introduce various issues, including bias or discrimination (Ferrara, 2023), security threats (Majdinasab et al., 2024a), and compromise of private information (Gupta et al., 2023). GenAI can also pose legal risks related to intellectual property concerns, including copyright infringement (Generative Artificial Intelligence and Copyright Law, [n. d.]). Many legal questions regarding the use of GenAI tools remain unanswered, and any answers will likely vary among jurisdictions and across cases.
@@
At the time of writing, there are several pending lawsuits against high-profile providers of GenAI tools including OpenAI (Authors Guild v. OpenAI Inc., 2023), StabilityAI (get, 2023), and Midjourney (Andersen v. Stability AI Ltd., [n. d.]), as well as active consideration by the U.S. Copyright Office and other governmental entities around the world of the copyright issues raised by the use of GenAI (Office, 2024a, 2023b; Council of the European Union, 2024b; Office, 2025). These matters involve several considerations, including the extent to which works associated with the use of GenAI, including prompts and models, are protected by copyright and, if so, who owns the rights to such works; whether the use of open-source software and other protected works as training data results in license violations or otherwise constitutes copyright infringement; and whether output resulting from the use of GenAI that is similar to preexisting work constitutes copyright infringement.
@@
Because any regulation of this space—whether by governmental entities or corporate entities—should take account of current and future practices, we aim in this article to identify developers’ views on (1) using GenAI tools for software development tasks and (2) associated legal concerns, focusing specifically on copyright law. This work aligns with requests from the U.S. Copyright Office seeking the opinions of experts, organizations, and individuals (e.g., developers) for active consideration in the development of future regulation (Office, 2023a).
@@
This article presents a study conducted by a joint team of SE and legal researchers that surveyed 574 software developers worldwide who reported using GenAI tools for software development tasks (particularly code generation). Through an online survey and follow-up interviews, we investigated developers’ opinions on potential emerging legal issues, the perception of what is copyrightable, ownership of generated code, and related considerations. We also sought to understand potential developer misconceptions, learn about the impact of GenAI on their work, and assess their awareness of licensing/copyright risks.
@@
Using qualitative and quantitative methods to analyze survey responses, we found that developer opinions on copyright issues vary broadly, particularly on the topic of model output ownership, and that many developers are aware of the nuances and complications associated with answering these complex legal questions. We discuss the results of our study, interpret the findings against the backdrop of U.S. law, and offer insights for future work. We focus on developers rather than other stakeholders, such as compliance officers and policymakers, because developers are often the ones who directly interact with GenAI during software development. We survey an international population of developers because software development is a global activity, particularly for open-source projects. In addition, large technology companies regularly employ individuals from various countries (and jurisdictions) around the world. Given this, it is appropriate to consider the thoughts, opinions, and perspectives of developers regardless of their geographical location.
@@
We make available our survey, results, and other artifacts for transparency and validation in an online replication package (ano, 2025).
@@ 2. Background and Related Work
Our work is intersectional, based on concepts drawn both from the legal domain and from software engineering. We discuss these concepts in this section as well as prior work related to our study.
@@ 2.1. Legal background
Unless it is in the public domain, computer code is generally protected by copyright in the U.S. so long as it constitutes an original work of authorship and is fixed in a tangible medium of expression, such as on a server (17 U.S.C. § 102, a). Similar provisions exist in the intellectual property legislation of other countries, including those in the European Union (Commission, 2024). The owner of a copyright—which may be one or more individuals or corporate entities—has several exclusive rights under U.S. copyright law, including the right to reproduce the work in copies, to create derivative works (works based on the copyrighted work), and to distribute the work, which also includes the right to authorize others to engage in such activities (17 U.S.C. § 106, [n. d.]). In addition, U.S. copyright law prohibits the removal or alteration of “copyright management information” defined as including “the name of, and other identifying information about, the author of a work,” although the limits of this provision and its application to GenAI are as yet uncertain (17 U.S.C. § 1202, b).
The use of GenAI implicates these rights in several ways that are currently under consideration by the U.S. Copyright Office and the courts. (Office, 2023b; The New York Times Company v. Microsoft Corporation, [n. d.]; Brittain, 2023; cop, 2023). For example, the plaintiffs in Doe v. GitHub (DOE 1 et al v. GitHub, [n. d.]), filed in November 2022, allege that Copilot removes copyright management information and violates the copyright licenses of the plaintiffs’ open source software by using that code as training data and generating code that is a near-identical reproduction of the plaintiffs’ code but without adhering to the terms of the associated licenses.
@@
Legislation both within and outside the U.S. also has implications for GenAI and copyright. Various U.S. states have attempted to enact legislation to govern GenAI (colorado-bill, 2024; california-regulation, 2024). In March 2024, the European Parliament approved the EU AI Act (Council of the European Union, 2024b), which, among other things, requires the providers of general-purpose AI models to “put in place a policy to comply with Union law on copyright and related rights, and in particular to identify and comply with, including through state-of-the-art technologies, a reservation of rights expressed pursuant to Article 4(3) of Directive (EU) 2019/790” (Council of the European Union, 2024a). This requires that creators of general-purpose AI models respect upstream copyrights—use due diligence in using copyrighted materials as part of their training data, obtain consent from copyright owners, and allow copyright holders to opt out—and respect downstream copyrights, mitigating the risk—often due to overfitting—of generating verbatim content that is too similar to the (copyrighted) original. In 2023, an early announcement of the EU Global Principles for Artificial Intelligence, which led to the EU AI Act, encouraged some organizations, including French media organizations France Médias Monde and TF1, to forbid companies such as OpenAI from mining their data (Karathanasis, 2023); that year also saw a temporary ban on ChatGPT in Italy (McCallum, 2023). Even countries like Japan, which initially affirmed that training AI models on protected content was not a violation of copyright law (Lanz, [n. d.]), have started to consider additional restrictions designed to protect copyright holders (Shinbun, 2023). Gervais et al. (Gervais et al., 2024) provide an in-depth overview of the then-current copyright landscape around the globe.
@@ 2.2. Technical background
Prior work relevant to GenAI and copyright issues focuses on research that may help answer some of the technological questions raised above. For example, there are techniques to potentially reveal whether a machine learning model was trained on a certain data record, such as membership inference attacks (Duan et al., 2024; Hu et al., 2022; Majdinasab et al., 2024b; Zhang and Li, 2023; Z. Yang et al., 2023). Such attacks can extract training data from the output of the target model, including private and personal information (Carlini et al., 2021). Memorization of training data is a known problem, including in models trained on code (Yang et al., 2024), with research suggesting that typical measures aimed at preventing verbatim extraction of training data are not sufficient to prevent the data from being obtained with other methods, such as by using style-transfer prompts (Ippolito et al., 2023). Tools such as CopyrightCatcher (CopyrightCatcher, [n. d.]) have been developed to detect the presence of copyrighted material in the output of models, demonstrating that this is an issue of concern for developers.
@@ 2.3. Related work
Prior research has explored several questions and concerns related to copyright and GenAI (Lucchi, 2023; Guadamuz, 2024). For example, Lee et al. (Lee et al., 2023a) outline the intricate GenAI supply chain and its intersection at various points with aspects of U.S. copyright law, revealing that answers to copyright questions may depend on the specific details of how an individual model was trained. Craig (Craig, 2024) has cautioned against applying or expanding copyright principles to address issues raised by GenAI without first carefully considering the unintended consequences of such an approach. Kugler (Kugler, 2024) provides a brief overview of the copyright challenges posed by GenAI and how various governments around the world have begun to address them.
Other research has considered the copyright implications of the emerging base model and the subsequent fine-tuning paradigm (Yew, 2024) and has studied the applicability of fair use for foundation models (Henderson et al., 2023). Additional research (Neel and Chang, 2023) presents a review of the discourse surrounding copyright and privacy issues related to LLMs. The research most closely related to our study was conducted by Liang et al. (Liang et al., 2024), who investigated why developers choose to use AI programming assistants and the challenges they face.
Our work builds on this related work by contributing an in-depth exploration of developers’ perceptions of the copyright issues surrounding the use of AI-based code-generation tools, both in the U.S. and elsewhere around the world. Understanding these perspectives allows future regulation in this area to be informed by beliefs on the ground rather than in a vacuum. The U.S. Copyright Office, to “inform the Office’s study and help assess whether legislative or regulatory steps in this area are warranted,” has actively sought public comment on copyright issues pertaining to artificial intelligence (Office, 2023a). Even after receiving over 10,000 comments, the Copyright Office continues to study this developing area (Office, 2024b, 2025). In the same spirit, our work aims to capture the practices and views of developers to provide a resource for policymakers.
@@ 3. Study Methodology
The purpose of this study is to investigate developers’ perceptions about licensing and copyright issues related to the use of GenAI technology for software development. (Although we use the term “software development” throughout the article, our study focuses on the use of GenAI for code generation as the main task that can lead to copyright and licensing issues.) The primary quality focus is to provide lawmakers, regulators, and other interested parties with a summary and analysis of relevant developer perspectives related to the copyright and licensing questions that arise when developing software with the help of GenAI. The context consists of 574 developers of open-source projects hosted on GitHub, whom we surveyed via an online survey and follow-up interviews.
@@
    What other legal concerns do developers anticipate as the use of GenAI increases? There are many other potential legal issues to consider as GenAI becomes more widely adopted, including data privacy, malicious content generation, and tort liability. In this RQ, we report developers’ perspectives on these topics, which extend beyond copyright law.
@@ 3.1. Survey design
To design our survey, we followed general guidelines (Groves et al., 2009) and SE-specific best practices (Pfleeger and Kitchenham, 2001; Kitchenham and Pfleeger, 2002a, b, c, d, 2003). The final questionnaire went through multiple iterations of review and improvements from six SE researchers and one law researcher, all authors of this article. All researchers contributed to the design of the study, but, in particular, the SE researchers contributed domain experience to the formulation of questions about procedures and practices relevant to developers, and the law researcher facilitated the identification of key areas that could be of concern related to copyright and licensing. We carefully formulated the survey questions and ensured that they were written clearly and concisely to avoid biasing the respondents. We also conducted a pilot study with graduate students from our research lab who actively use GenAI tools. Based on their feedback, we further refined our questionnaire, improving questions that were ambiguous or difficult to understand.
@@
  • The section on “Current use of GenAI for coding” (12 questions) asked about respondents’ experience using GenAI tools for code generation (C1, C2, C8), including the tools they used (C3); the development tasks for which they used the tools (C4); the benefits and challenges of using such tools (C5, C6); whether they read the Terms of Service for tools (C7); organizational policies for using such tools (C9, C12); and procedures to document tool usage (C10, C11).
  • The section on “Understanding and Perception of Copyright Issues” (14 questions) asked about developers’ perceptions of copyright issues surrounding GenAI, including their views on if and how copyright, litigation, and legal risks factor into the use of GenAI (U1, U2, U3); the effects of these concerns on developers’ workflows (U4); whether the use of source code in training models should require monetary compensation and/or attribution (U5, U6); the use of code in model training without permission (U7); having models produce code similar to code they created (U8); reusing without permission a prompt that they created (U9); ownership and copyrightability of generated code (U11, U12, U13); and other, broader concerns about copyright/legal issues stemming from development tasks with AI-based tools (U14).
  • The “Demographics” section (8 questions) asked about respondents’ age (D1); experience with SE and AI (D2, D3); their primary programming languages (D4); whether they develop open source or proprietary software (D5); their location (D6); and if they have received training on copyright law (D6, D7).
The survey was designed to be completed in about 15 to 20 minutes and consisted mainly of multiple-choice questions (26), complemented by open-ended questions (7) that allowed participants to elaborate on their responses. Because we included logic in our survey to ask only those questions relevant to a participant’s indicated experience, not every participant saw every question. For example, if a participant indicated that they were not familiar with the practice of generating source code using GenAI (C1), we did not ask them questions about generating source code (C2-C12). The complete survey text and a brief description are found in our replication package (ano, 2025).